The odds that a book will be favourably reviewed by three independent critics are 5 to 2, 4 to 3 and 3 to 4 respectively.what is the probability that of three reviews a majority will be favourable.

Grade:12th Pass

1 Answers

Harsh Patodia IIT Roorkee
8 years ago
R1, R2 and R3 be probability of favourable reviews
R1 = 5/7
R2 = 4/7
R3 = 3/7
For Majority atleast 2 should be favourable.
i.e. R1xR2xR3’ +R1x R2’xR3 + R1’xR2xR3 + R1R2R3 ( where R’ represent not favored)
Just plugin values and get the answer.